### v3

# Knowledge Bridger: Towards Training-Free Missing Modality Completion

Abstract
Previous successful approaches to missing modality completion rely on carefully designed fusion techniques and extensive pre-training on complete data, which can limit their generalizability in out-of-domain (OOD) scenarios. In this study, we pose a new challenge: can we develop a missing modality completion model that is both resource-efficient and robust to OOD generalization? To address this, we present a training-free framework for missing modality completion that leverages large multimodal model (LMM). Our approach, termed the “Knowledge Bridger”, is modality-agnostic and integrates generation and ranking of missing modalities. By defining domain-specific priors, our method automatically extracts structured information from available modalities to construct knowledge graphs. These extracted graphs connect the missing modality generation and ranking modules through the LMM, resulting in high-quality imputations of missing modalities. Experimental results across both general and medical domains show that our approach consistently outperforms competing methods, including in OOD generalization. Additionally, our knowledge-driven generation and ranking techniques demonstrate superiority over variants that directly employ LMMs for generation and ranking, offering insights that may be valuable for applications in other domains.
1 Introduction
Multimodal learning [39, 21] is a foundational task in artificial intelligence that enables models to integrate diverse data types—such as text, images, and audio—to enhance their representational and reasoning capabilities [47, 34, 33]. However, real-world applications often encounter incomplete or missing modalities due to factors like noise, sensor failures, or privacy constraints, which can compromise a model’s robustness and generalizability. Missing modality completion (MMC) addresses this issue by imputing or reconstructing absent modalities based on available data, allowing models to fully leverage multimodal information. Recent MMC methods [17, 56, 26, 4] have demonstrated promising results across various applications. For instance, in medical diagnostics [63, 7], MMC models can reconstruct absent diagnostic data (e.g., MRI, CT, and X-rays) by leveraging existing medical reports or related images.
Despite these advancements, MMC methods often have limited out-of-domain (OOD) transferability, typically requiring retraining on new domain data to maintain performance. For example, some MMC models [65, 36] developed for sentiment analysis need substantial adaptations to be effective in applications like medical imaging or autonomous driving, leading to significant human and computational costs. To address these challenges, recent works have proposed domain-agnostic solutions, such as missing modality tags [61] to aid in predicting absent modalities or prompt-learning techniques [17, 26] that dynamically adjust fusion strategies. While these methods lower domain adaptation costs, they still require extensive training data, limiting their effectiveness in data-scarce domains, such as rare disease analysis. This raises a crucial question: Can we develop an MMC model that achieves both low-resource dependency (e.g., minimal computational and domain-specific data requirements) and strong OOD capability?
Recent advancements [2, 32, 51, 20] in large multimodal models have highlighted their strong OOD capabilities and adaptability to new tasks with minimal resources [24]. In this work, we explore the potential of leveraging LMM to effectively address the MMC challenge. Before progressing, it is essential to distinguish between modality generation and missing modality completion. Modality generation typically focuses on creating new modalities from available information, prioritizing diversity and creativity in the generated content. In contrast, missing modality completion emphasizes accuracy over diversity, reconstructing absent modalities to maintain semantic coherence and improve task performance.
Addressing MMC with LMM presents two primary challenges: 1) generation: applying constraints on modality generation to ensure fidelity to the missing content, and 2) ranking: selecting the most appropriate completion from generated candidates. Effective use of LMM for these tasks requires embedding sufficient domain knowledge to guide accurate interpretation and reconstruction of missing content. For example, generating an accurate image in a vision-language task can be challenging with only brief textual descriptions; however, incorporating detailed descriptions of entities and their interactions allows LMM to more accurately reconstruct missing data. Additionally, such prior knowledge aids LMM in ranking candidate completions by identifying the most semantically plausible options. Leveraging LMM’s in-context learning capability [59], we can non-intrusively embed prior knowledge, such as using the Chain-of-Thought (CoT) approach [60], to enhance the model’s understanding of missing modalities.
Based on these insights, we propose a novel, training-free MMC approach, termed “Knowledge Bridger”, which autonomously mines multimodal knowledge, generates missing modalities, and ranks the best completions. This method comprises three main modules: a knowledge modeling module, a knowledge-driven modality generation module, and a ranking module.
In the knowledge modeling stage, we employ LMM to analyze available modalities and extract key elements such as objects, interactions, and attributes using the CoT approach. For specialized fields like medical imaging, we use in-context learning to integrate domain-specific knowledge, reducing reliance on extensive target domain data. Following knowledge extraction, we construct a knowledge graph to represent the relationships and attributes of available and missing modalities, providing LMM with a structured reference for generating missing content. To ensure high-quality completion, the knowledge-driven generation module utilizes this graph to guide LMM in generating specific content with precise details, such as object locations and observable attributes. Finally, the ranking module, informed by expert knowledge, assesses each generated completion by computing graph and representational similarity scores. Specifically, a graph similarity score is derived from comparing the knowledge graphs of the available and generated modalities, while a representational similarity score is calculated using models like BLIP [28] and CLIP [45]. A weighted average of these scores offers the final assessment.
Our extensive experiments indicate that our method markedly improves MMC performance in both general and OOD scenarios. We also find that our approach scales effectively with LMM, with larger models yielding higher-quality completions. For instance, using OpenAI’s GPT-4o [20] results in a significant performance boost across metrics compared to models with 72B or 7B parameters. The visualization results show that our method outperforms the conditional generation variant in generating missing modalities. Our contributions are summarized as follows:
We introduce a training-free pipeline to address missing modality completion, leveraging LMM to automatically extract multimodal knowledge, generate missing modalities, and rank completions. To our knowledge, this is the first work to apply LMM to MMC tasks.
We delve into a modality-agnostic unified strategy for both missing modality completion and ranking. This approach allows us to focus on defining domain-specific knowledge without the necessity of intricate fusion methods or implementing a specialized training pipeline.
We present extensive experimental evidence demonstrating that our method facilitates domain transfer, outperforming other MMC methods in both general and OOD scenarios. Additionally, our generated completion data improves the performance of other MMC models.
2 Method
2.1 Overview
Our approach aims to construct a training-free MMC pipeline by leveraging a pre-trained LMM. This pipeline extracts and models knowledge from available modalities, and subsequently uses this knowledge to generate missing modality data and select the most appropriate generated candidate. The pipeline is depicted in Fig. 1 and consists of three steps:
Step 1: construction of a knowledge graph from available modalities (in section 2.2). The goal of this step is to utilize the general knowledge a pre-trained LMM to understand the content of each modality and their interrelationships.
Step 2: knowledge-driven generation (in section 2.3). In this step, the LMM employs the knowledge graph to extract specific details about the missing modality, including the number and attributes of objects. A corresponding modality generator then uses this information to generate the required missing modality.
Step 3: knowledge-based ranking (in section 2.4). This step aims to compute a quality score for the generated missing data by assessing the graph similarity and semantic similarity between the missing and available modalities.
2.2 Knowledge Graph Modeling
Our objective is to automatically extract knowledge from available modalities to support the generation and ranking of any missing modalities. In our context, ‘knowledge’ refers to information encapsulating the characteristics of existing modalities, enabling the generation and ranking modules to create semantically consistent missing data. Extracting relevant knowledge from unknown domains, however, presents significant challenges. From a knowledge graph perspective, constructing a meaningful, modality-specific graph requires predefined nodes and relationships. In a training-free context, predefining these elements becomes particularly challenging.
To overcome this, we develop an automatic entity and relationship mining strategy using an LMM. This strategy leverages the extensive prior knowledge and OOD capabilities of LMM to identify entities and relationships across various modalities, even without predefined elements. Recent study [24] highlight LMM’s potential for zero-shot learning and reasoning.
Building on this analysis, LMM can be utilized to extract elements from available modalities using prompts. To enhance scalability, we propose the following extraction rule: {Entity: Reasoning Prompts}. For instance, to identify potential objects, we could use: {‘Objects’: ‘Identify the major objects in the [modality-type].’}. This rule allows us to incorporate object relationships and interaction data111The complete rules and prompts are provided in the appendix.. To improve adaptability across domains, domain-specific prior knowledge—such as histological and clinical diagnostic information for medical image analysis—can be included. This approach offers two main advantages: 1) reducing misconceptions when LMM operates in a new domain, and 2) enhancing its reasoning capabilities for novel entities.
LMM can further consolidate the extracted information into a modality-specific knowledge graph. A straightforward approach is to guide LMM to extract potential entity-relation pairs from the collected data. However, this method is constrained by the context window length, and an excessive number of extraction rules may lead to overlooked entity-relation pairs. To mitigate these limitations, we implement the Chain-of-Thought (CoT) method. Specifically, LMM is first directed to produce concise responses for each rule, followed by extracting unique entity-relation pairs from these responses. This stepwise decomposition improves both the accuracy of responses for each rule and the synthesis of information. Importantly, we extract and retain data solely from current modalities to avoid interference from unrelated information, thereby enhancing LMM’s reasoning efficiency.
2.3 Knowledge-driven Generation
The goal of missing multimodal generation is to understand the content in available modalities and generate missing ones that align semantically. Two critical factors impact the quality of the generated missing modalities: understanding multimodal content and maintaining consistency. We previously discussed using LMM for content comprehension and knowledge graph extraction. Here, we explore using LMM for ensuring consistency and guiding generation. For convenience, we use image-text pairs as our study objects. For example, when an image is available, we aim to generate text that closely matches real data. A basic approach is using LMMs to describe the image directly. However, this method involves significant randomness. First, the form of the missing text is unknown—it could be a caption, summary, or description. Second, we cannot precisely specify the subject of the missing text.
To solve this, we propose a knowledge-driven entity alternation strategy. Using domain knowledge and extracted knowledge graphs, we select related entities.For instance, if missing data focuses on an entity such as an ‘objection,’ we traverse elements within the knowledge graph related to ‘objection.’ We then adopt a multi-view generation manner, allowing the LMM to generate missing information with each element as the subject, while encompassing all nodes and attributes within the knowledge graph. These outputs are stored as standardized text descriptions, reducing randomness, enhancing result retrieval, and offering better control and interpretability. With these descriptions, modality generators can create missing data. For missing images, entity-based descriptors can guide conditional diffusion methods. For missing text, LMM process these descriptions to generate outputs. This approach works across various fields, using mature generation models with domain knowledge to create needed data without extra training. However, relying only on this may not guarantee full accuracy, which we discuss in the next section.
2.4 Knowledge-based Ranking
To achieve automatic ranking of generated missing data based on the provided knowledge, we introduce graph similarity and representation similarity. Graph similarity is computed as the average cosine similarity score of the adjacency matrices of two graphs, as presented in the following equation:
where and represent the -th row vectors of adjacency matrices and , respectively. refers to the Euclidean norm of the -th row: . and represent the rows and columns of the adjacency matrix, respectively. This metric reflects the degree of similarity between the two graphs, and its value is normalized between 0 and 100. On the other hand, we compute the representation similarity between the generated and available modalities to reflect semantic consistency. For consistency, we similarly employ cosine similarity to compute the similarity between two representations, expressed as , where and are the vectors of two modalities. Inspired by [3], we directly utilize CLIP [45] and BLIP [28] to obtain semantic embeddings for each modality. Finally, we derive the following equation to compute the generation quality score between any pair of available and missing modalities:
where and represent available and missing modalities, respectively. The functions , , and are used to obtain the adjacency matrix, CLIP’s embedding, and BLIP’s embedding of the given modality, respectively. We argue that can comprehensively assess two critical factors: knowledge structure similarity and semantic consistency. A higher value indicates a higher quality of the generated missing modality. Our method ultimately outputs the generated missing modality with the highest score.
3 Experiments
3.1 Setup
Baselines. To validate the effectiveness of our proposed method, we selected several baseline methods, categorized into two types: imputation-based and non-imputation-based. Imputation-based methods aim to restore missing modalities by learning relationships among the representations within each modality; for this category, we choose MMIN [65] and DiCMoR [56] as baselines. In contrast, non-imputation-based methods bypass the need to restore missing modalities or predict their representations, allowing downstream tasks to be completed without these steps. We choose MPLMM [17] and MPMM [26] as baselines for this category. Additionally, we introduce a simple baseline that completely removing missing data. This baseline utilizes the pre-trained CLIP model as multimodal backbones. Subsequently, the modality features are concatenated and fed into a classification head (a single-layer MLP) to ultimately obtain the probability for each category.
Dataset. To evaluate the domain transferability of our method, we employ two general domain datasets and one OOD dataset. The general domain datasets include COCO-2014 [31] and MM-IMDb [1]. COCO-2014 [31] is a large-scale vision-language dataset containing approximately 81 training samples and 41 validation and test samples, with objects classified into 80 categories. In our setup, we treat it as a multimodal multi-label classification dataset, using the validation set as the test set. MM-IMDb [1] is a large-scale movie genre classification dataset with around 25,000 movies from IMDb, each represented by its movie poster and plot summary, and annotated with 27 multi-label genre tags. In our setup, we divide the dataset into 18,160 samples for the training set and 7,799 samples for the test set. On the other hand, we employ the IU X-ray [8] dataset as our OOD dataset. This dataset includes 6,674 training and 756 testing samples. We further expanded the original 14 observations into 105 finer-grained categories based on location and severity, creating a long-tail dataset.
Evaluation Metrics. For multimodal multi-label classification on three datasets, we report the performance of comparison methods using the macro F1-score and Average Precision (AP). We report the average result of five different random seeds. Additionally, we introduce the mean similarity score (denoted SS) to evaluate the generation quality of our method and other imputation-based methods. Specifically, we utilize the vision and text backbones of the pre-trained CLIP [45] to compute embeddings for all ground-truth and generated missing modalities. The cosine similarity between each pair of embeddings is then computed and averaged, with scores ranging from 0 to 100. A higher score indicates better generation quality.
Implementation Details. We employ Qwen2-VL-7B [55] as our default large multimodal model222More implementation details are presented in the appendix.. It is specifically designed to handle both visual and textual inputs, enabling it to process and interpret complex multimodal content. For image reconstruction, we apply Stable Diffusion XL (SDXL) 1.0 [43] as the restoration module for general domains. SDXL 1.0 is an advanced text-to-image diffusion model that can generate images according to a given prompt. Additionally, for the restoration of chest X-ray modality, we use Cheff [58], a cascaded chest X-ray latent diffusion model. By default, we generate 5 candidates for the missing modality during the generation process. Since the chosen LMM has already been pre-trained on the COCO dataset, to ensure a fair comparison, we employ pre-trained CLIP [45] vision and text encoders to replace the modality backbones of the comparison methods. We conduct all experiments on the PyTorch 2.4.0 [41] platform, running on Ubuntu 20.04 LTS utilizing 4 GPUs (NVIDIA GeForce RTX 4090 with 24 GB of memory). In our setting, we conduct the missing rate to simulate the missing modality scenario during training.
| COCO-2014 [31] | MM-IMDb [8] | |||||||||||||||||
| Missing Rate | 0.3 | 0.5 | 0.7 | 0.3 | 0.5 | 0.7 | ||||||||||||
| Method | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS |
| Baseline (complete) | F1: 78.3 AP: 84.6 SS: - | F1: 56.2 AP: 62.7 SS: - | ||||||||||||||||
| Baseline (remove missing) | 75.8 | 80.8 | - | 73.1 | 79.3 | - | 70.3 | 77.5 | - | 51.8 | 58.5 | - | 50.3 | 56.1 | - | 47.2 | 53.9 | - |
| MPMM [26] (CVPR’23) | 76.2 | 82.0 | - | 74.7 | 80.5 | - | 71.0 | 78.8 | - | 53.6 | 59.8 | - | 51.1 | 56.9 | - | 48.5 | 55.7 | - |
| MPLMM [17] (ACL’24) | 77.1 | 82.6 | - | 75.2 | 81.3 | - | 72.3 | 80.1 | - | 53.9 | 60.3 | - | 52.8 | 57.3 | - | 49.1 | 56.2 | - |
| MMIN [65] (ACL’21) | 73.2 | 78.3 | 37.1 | 71.4 | 77.5 | 36.7 | 70.5 | 76.4 | 36.0 | 50.1 | 53.8 | 26.7 | 49.5 | 51.7 | 26.3 | 44.6 | 50.8 | 24.4 |
| DiCMoR [56] (CVPR’23) | 65.3 | 74.4 | 34.3 | 59.7 | 67.1 | 33.5 | 55.3 | 64.0 | 31.9 | 49.2 | 54.7 | 26.9 | 43.7 | 50.8 | 25.9 | 30.5 | 41.7 | 23.1 |
| Ours (Qwen-VL-2B) | 76.1 | 81.8 | 39.8 | 74.3 | 79.9 | 37.1 | 70.7 | 78.2 | 36.3 | 52.4 | 58.9 | 28.8 | 51.9 | 57.1 | 28.3 | 50.3 | 56.5 | 28.1 |
| Ours (Qwen-VL-7B) | 77.5 | 82.9 | 40.4 | 77.5 | 82.8 | 40.2 | 77.9 | 83.5 | 38.2 | 54.7 | 60.9 | 33.5 | 54.9 | 61.3 | 32.7 | 55.2 | 61.8 | 32.3 |
| Complete Baseline | -2.8 | -1.7 | - | -0.8 | -1.8 | - | -0.4 | -1.1 | - | -1.5 | -1.8 | - | -1.3 | -1.4 | - | -1.0 | -0.9 | - |
| SOTA | +0.4 | +0.3 | +3.3 | +2.3 | +1.5 | +3.7 | +5.6 | +3.4 | +2.5 | +0.8 | +0.6 | +6.6 | +2.1 | +4.0 | +6.4 | +6.1 | +5.6 | +7.9 |
| Missing Rate | 0.3 | 0.7 | ||||
| Method | F1 | AP | SS | F1 | AP | SS |
| Baseline (complete) | F1: 57.0 AP: 75.7 SS: - | |||||
| Baseline (remove missing) | 49.1 | 71.4 | - | 31.5 | 56.2 | - |
| MPMM [26] (CVPR’23) | 49.9 | 71.8 | - | 36.8 | 61.4 | - |
| MPLMM [17] (ACL’24) | 49.3 | 72.7 | - | 35.2 | 61.9 | - |
| MMIN [65] (ACL’21) | 37.3 | 64.2 | 17.3 | 26.7 | 50.1 | 10.2 |
| DiCMoR [56] (CVPR’23) | 40.5 | 69.1 | 18.1 | 29.8 | 53.6 | 13.3 |
| Ours (Qwen-VL-2B) | 51.4 | 72.0 | 21.9 | 41.1 | 68.9 | 17.7 |
| Ours (Qwen-VL-7B) | 53.6 | 73.9 | 22.6 | 46.3 | 70.5 | 19.8 |
| Complete Baseline | -3.4 | -1.8 | - | -10.7 | -5.2 | - |
| SOTA | +3.7 | +1.2 | +4.5 | +9.5 | +8.6 | +6.5 |
3.2 Quantitative Analysis
In Tables 1 and 2, we present the results of our method compared to others across different missing data ratios, as measured by F1, AP, and mean similarity score (SS) on three datasets333More results (more modalities) are presented in the appendix.. To comprehensively evaluate the performance of various approaches, we introduce two baselines: one trained on complete data and another trained on data with missing entries removed. Additionally, to fair comparison regarding the scale of model parameters, the proposed method based on the Qwen-VL-2B were incorporated.
Results from general domain datasets (COCO-2014 and MM-IMDb) indicate that most MMC methods outperform the baseline where missing data is simply removed. Our method demonstrates superior performance across different missing rates, particularly at a missing rate of 0.7, where it shows significant improvements in both F1 and AP metrics. We attribute these enhancements to the synthetic data generated by our method, which appears to bolster downstream task performance. This phenomenon is also observed in [16, 54]. However, it should be noted that the synthetic data may lead to a slight decline in mean similarity score.
On the other hand, results from the OOD dataset highlight our method’s superior domain adaptation capability compared to other methods. Our approach not only excels in classification metrics but also significantly surpasses other imputation-based methods, such as MMIN and DiCMoR, in missing data generation metrics (SS). This advantage can be credited to our method’s ability to harness the few-shot learning and in-context learning capabilities of LMM, effectively mitigating the impact of varying degrees of data incompleteness.
3.3 Ablation Study
Q1: Does ‘knowledge’ really help MMC? We study the impact of different components under general and OOD scenarios with the same missing rate. As shown in Table 3, we report the results of various components. For the generation process, we analyze the variant without knowledge modeling444We directly use LMM to generate descriptions of available modalities. (row 1) and the variant with random ranking (row 2). Additionally, in examining the ranking module, we assess the variant with random ranking (row 3), the variant without knowledge graph ranking (row 4), and the variant without using semantic similarity score ranking (row 5). The results demonstrate that the variant without knowledge modeling shows a significant decline across all metrics under OOD scenarios. Furthermore, the semantic similarity ranking strategy in the ranking module is crucial for the effectiveness of the MMC. In general, the study reveals that knowledge modeling is the most critical component of our method.
| MM-IMDb | IU X-ray | ||||||
| Missing Rate | 0.7 | 0.7 | |||||
| Variants | F1 | AP | SS | F1 | AP | SS | |
| 0 | Baseline (Qwen-VL-7B) | 55.2 | 61.8 | 32.3 | 46.3 | 70.5 | 19.8 |
| 1 | w/o Knowledge Modeling | -1.3 | -3.6 | -8.8 | -17.5 | -29.2 | -13.7 |
| 2 | + Random Ranking | -1.6 | -4.1 | -9.9 | -19.3 | -31.8 | -15.0 |
| 3 | Random Ranking | -0.5 | -2.7 | -0.6 | -3.8 | -7.1 | -4.7 |
| 4 | w/o Knowledge Ranking | -0.2 | -0.8 | -0.1 | -1.9 | -2.7 | -1.1 |
| 5 | w/o Semantic Ranking | -0.2 | -1.0 | -0.3 | -2.4 | -3.3 | -1.6 |
Q2: Does the scale of model parameters affect MMC? We evaluate the performance of our approach using different model scales under at various missing rates. Specifically, we conduct the 2B, 7B, and 72B variants of the open-source large model Qwen-VL [55], as well as OpenAI’s GPT-4o [20] on the IU X-ray dataset. For a fair comparison, we leverage these models solely for handling the knowledge modeling and integration in steps 1 and 2 of our method, keeping the modality generators unchanged. As illustrated in Fig. 2, GPT-4o demonstrates significant superiority and robustness. These results demonstrate that as the scale of model parameters increases, the quality of knowledge modeling in our approach also improves.
Q3: Does generating more missing data improve performance? We evaluate the impact of different numbers of generation candidates in step 2 of our method on the MM-IMDb and IU X-ray datasets. Specifically, we conduct the settings with generation candidates , and the results are presented in Table 4. The findings indicate that generating only one candidate significantly reduces the missing generation quality. While using a higher number of generation candidates can lead to slight improvements across all metrics, it also substantially increases inference time. Therefore, setting provides a balance between inference time and performance.
| MM-IMDb | IU X-ray | ||||||
| Missing Rate | 0.7 | 0.7 | |||||
| Candidates | time | F1 | AP | SS | F1 | AP | SS |
| (default) | 40s | 55.2 | 61.8 | 32.3 | 46.3 | 70.5 | 19.8 |
| 7s | -0.7 | -2.8 | -7.3 | -8.5 | -16.4 | -14.7 | |
| 83s | +0.4 | +0.9 | +0.1 | +0.0 | +0.1 | +0.0 | |
| 132s | +0.4 | +0.9 | +0.1 | +0.1 | +1.3 | +0.3 | |
3.4 Visualization Analysis
To better understanding the differences between our approach and conditional generation, we present completion results produced by our method, as shown in Fig. 3. We provide visualization results across different missing modalities for three datasets555More visualization are presented in appendix.. In the general domain, our knowledge modeling module focuses on understanding the quantity of objects, their attributes, and the contextual environment. The results in Fig. 3a and Fig. 3b indicate that our method is more similar with the original missing modality compared to direct generation approaches.
In the medical domain, incorporating knowledge of different lesions enables the LMM to understand the relationships between various regions in chest X-rays and the content described by the modality. Fig. 3c and Fig. 3d show that our method provides a more reliable strategy for missing data completion than direct generation.
4 Related Work
4.1 Missing Multimodal Learning
Missing multimodal learning enables models to train and inference effectively even when some modalities are absent. The missing multimodal learning methods can be categorized into imputation-based and non-imputation-based approaches. Imputation-based methods predict or recover missing data by learning interactions between modalities. Typical strategies include filling with special values such as zero [40] or the average value [62], which are unsuitable for high-dimensional data. Advanced methods like SMIL [36] use meta-learning to simulate missing modalities. Wang et al. [53] propose a multitask framework for missing modality learning, which learns shared features across different tasks and uses these shared features to reconstruct missing modalities. Lian et al. [30] use graph neural networks to explore relationships between modalities and recover missing ones, while Zhao et al. [65] and Pham et al. [42] leverage cycle consistency to capture cross-modal interactions for predicting missing modalities.
On the other hand, non-imputation-based methods focus on fusion strategies or missing indicators instead of predicting missing data [22]. Wang et al. [57] and Han et al. [19] use translational semantics for implicit fusion. Ma et al. [37] leverage transformers’ variable input capability. Lee et al. [26] and Guo et al. [17] introduce learnable missing semantic tokens, while Zeng et al. [61] use fixed tags to identify missing inputs. Current methods often require pre-training with complete data and show limited domain transfer. Our approach leverages LMM to predict missing modalities using domain knowledge without extra pre-training, employing few-shot learning for flexible adaptation.
4.2 Large Multimodal Models
Large multimodal models (LMMs) integrate diverse data types, such as text, images, and audio, to enhance understanding and interaction capabilities. Early works focused on fusion techniques to combine these modalities, like early [5], late [12], or hybrid fusion [44]. Recent advancements [51, 52], such as OpenAI’s GPT-4o [20] and Google’s PaLM-E [9], have achieved remarkable success by increasing the scale of both pre-training data and model parameters. These private ultra-models (greater than 100 billion parameters) exhibit robust generalization across various domains, which enables applications in image captioning, visual question answering, and multimodal dialogue. On the other hand, some open-source, medium- to small-scale (less than 100 billion parameters) models, such as LLaVA [32], Qwen-VL [2], and InternVL [6], have matched the performance of these ultra-large models. The common feature among them is their rich prior knowledge, understanding, and OOD capabilities, which form the foundation of training-free missing multi-modality completion.
4.3 Conditional Generation
Conditional generation involves the creation of new data samples based on specific input data or attributes, allowing models to produce outputs tailored to given prompts or features. Early foundational work in this area includes Generative Adversarial Networks (GANs) [15] and Variational Autoencoders (VAEs) [23], with notable advancements such as conditional GANs [38] and Conditional VAEs (CVAEs) [49]. Recent progress has introduced diffusion models [50, 29, 48], which have demonstrated the ability to generate more realistic content, as exemplified by systems like DALL-E [46] and Stable Diffusion [43]. Zhang et al. introduce ControlNet [64], which significantly enhances user control over generated content by utilizing detailed, structured prompts. The key distinction between these approaches and our method lies in their respective objectives. While the former focuses on generating diverse content in response to prompts, Our method goes beyond this goal by requiring the model to understand interactions between modalities to accurately complete missing content.
5 Discussion and Limitations
Why we need the training-free MMC? Previous approaches for handling missing modalities have demonstrated limited generalizability across different domains and constrained data flexibility. For example, recent prompt-based methods such as MPMM [26] and MPLMM [17] have shown significantly lower performance in general domains compared to their effectiveness in the medical domain, as reflected in our experimental results. Additionally, these methods depend on pre-trained models that are highly tailored to the target domain. In data-hungry scenarios, such as the medical field where data often contain missing elements, these constraints greatly impact the reliability of model decisions. In recent years, LMMs have exhibited remarkable zero-shot capabilities across various domains. Consequently, leveraging these models in a training-free manner to address the MMC problem presents a promising and cost-effective research direction. Our experimental results corroborate this perspective, and we hope these findings will encourage the community to further explore the potential of LMMs for addressing MMC challenges.
Why ‘knowledge bridger’ can help MMC? We believe that the key to addressing the MMC problem using LMM lies in accurately generating missing data and effectively ranking the generation candidates. Our experimental results show that directly employing LMM to generate the missing modality does not ensure the accuracy required for MMC. Our proposed approach first involves mining internal knowledge from the available modalities, enabling LMM to understand intra-modal interactions. This structured knowledge is then used to address challenges in both generation and ranking. However, using LMM inherently leads to hallucinations. In future work, we plan to employ more robust knowledge extraction methods, such as retrieval-augmented generation [18], to mitigate hallucinations666Potential hallucinations of our method are presented in appendix. and further enhance the accuracy of missing data generation.
Limitations. Our method focuses exclusively on image and text modalities, leaving its performance on other modalities, such as speech and depth, yet to be explored. The approach emphasizes the automatic extraction of inter-modal knowledge and the completion of missing modalities through domain knowledge. Thus, in the future, adaptation to other modalities is possible by defining a more comprehensive modality knowledge. Some promising works [14, 35] show that one modality, such as image or text, can be connected to any other modality. Additionally, we observe that while our method enhances classification performance under a high missing rate (e.g., 0.7), it paradoxically results in a decrease in the similarity scores of the completed modalities. Therefore, there remains substantial potential for further exploration to develop more robust generation and ranking strategies in the future.
6 Conclusion
In this paper, we propose a novel MMC challenge and introduce a training-free MMC pipeline called ‘Knowledge Bridger’ based on LMM. Our findings demonstrate that LMM can autonomously mine knowledge across modalities and leverage this structured knowledge to enhance the accuracy of missing modality generation and ranking. Furthermore, experimental results indicate that as the model’s parameter scale increases, our method shows improved knowledge extraction capabilities.
Acknowledgment. This work is supported by the Guangdong Natural Science Funds for Distinguished Young Scholars (Grant 2023B1515020097), the AI Singapore Programme under the National Research Foundation Singapore (Grant AISG3-GV-2023-011), the Beijing Natural Science Foundation (No.4234086), the Natural Science Foundation of China (No. 62192782), and the Lee Kong Chian Fellowships.
Supplementary Material
G Prior Knowledge Rules and Prompts
We report the knowledge extraction rules used in our method and the prompts used in the knowledge-driven generation respectively.
G.1 Prior Knowledge
To simplify the research problem, we utilize only a set of basic instructions to direct the large language model (LMM) focus toward the desired multimodal content extraction. For the general domain, we predefine the extracted knowledge to include major objects, the quantity of each object, and their corresponding attributes and styles. Given the strong reasoning capabilities of LMM in the general domain, it is not necessary to differentiate between input modalities. However, to mitigate hallucinations generated by the LMM, we limit the number of major objects. In our experiments, extracting between 5 to 7 objects show optimal.
In the medical domain, it is necessary to pay attention to the distinctions between different modalities. This is because LMM may not inherently comprehend the knowledge required for out-of-domain (OOD) scenarios, necessitating clear identification of the modality being processed and the content to be understood. In our approach, for X-rays, we instruct the LMM to focus on anatomical structures, clinical significance, abnormal findings, and report generation.
For reports, in addition to the information extracted from X-rays, we direct the LMM to consider the locations of the anatomical structures mentioned and any additional characteristic details present in the report.
G.2 Knowledge Extraction with Chain-of-Thought
We employ an LMM with Chain-of-Thought (CoT) [60] reasoning to extract knowledge according to the aforementioned rules. This approach helps reduce the computational strain associated with long-problem reasoning and enhances the overall accuracy of problem-solving. For the general domain, our final instruction prompt is designed as follows:
For medical domain, we have:
Next, we require the LMM to integrate the CoT results according to the following instructions:
We strictly require the LMM to return structured information in the following format:
After extracting the aforementioned structured information, we employ LMM to transform these knowledge into the form of a knowledge graph. To simplify the process, we represent relationships on the graph using a triplet structure (nodes and edges):
G.3 Knowledge-driven Generation
We employ the knowledge graphs extracted by the aforementioned process as input and employ the LMM to process this information to generate meaningful descriptions of missing modalities. Various modality generators are then utilized to produce the missing information based on these descriptions. For missing image, predictions are based on observable text:
For missing text, the same method is applied to generate descriptions of the missing content, which are then refined by the LMM to produce the required missing text.
G.4 Knowledge-based Ranking Pseudo-code
The pseudo-code for the ranking process is shown in Alg. 1.
H Implementation Details
GPUs Details. We conduct all experiments on the PyTorch 2.4.0 [41] platform, running on Ubuntu 20.04 LTS utilizing 4 GPUs (NVIDIA GeForce RTX 4090 with 24 GB of memory).
Deploy Efficient Large Multimodal Model. We deploy the Qwen-VL[55] large model using vLLM [25]. vLLM is an LLM serving system that achieves (1) near-zero waste in KV cache memory and (2) flexible sharing of KV cache within and across requests to further minimize memory usage. We deploy versions with 2B777https://huggingface.co/Qwen/Qwen2-VL-2B-Instruct, 7B888https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct, and 72B parameters. Specifically, due to hardware constraints, we utilize the 72B quantized version with Int-8 precision available on Hugging Face: Qwen2-VL-72B-Instruct-GPTQ-Int8999https://huggingface.co/Qwen/Qwen2-VL-72B-Instruct-GPTQ-Int8. For all versions, we maintain an 8K context window length and support a maximum of four image queries. For each query, we set the maximum number of tokens to 512 and use a temperature of 0.1.
Generators Settings. For image reconstruction, we apply Stable Diffusion XL (SDXL) 1.0 [43] as the restoration module for general domains. SDXL 1.0 is an advanced text-to-image diffusion model that can generate images according to a given prompt. Additionally, for the restoration of chest X-ray modality, we use Cheff [58], a cascaded chest X-ray latent diffusion model. By default, we generate 5 candidates for the missing modality during the generation process.
Missing Modality Simulation In our setting, we conduct the missing rate to simulate the missing modality scenario during training. Specifically, we calculate the number of missing samples in different datasets under a given missing rate and then randomly mark the text and image modalities of these samples as missing with a probability of 0.5. To ensure the reproducibility of the experiment, we perform multiple simulations using the same set of missing samples. Finally, we retrain the baseline model, which was initially trained on complete modalities, using the data with imputed missing modality and report the performance across various metrics.
I More Experimental Results
I.1 Table of Quantitative Results
We present additional quantitative analysis results, as shown in Table 5.
| Missing Rate | 0.5 | ||
| Method | F1 | AP | SS |
| Baseline (complete) | F1: 57.0 AP: 75.7 SS: - | ||
| Baseline (remove missing) | 42.1 | 63.9 | - |
| MPMM [26] (CVPR’23) | 42.8 | 64.2 | - |
| MPLMM [17] (ACL’24) | 43.3 | 65.9 | - |
| MMIN [65] (ACL’21) | 34.8 | 60.1 | 17.0 |
| DiCMoR [56] (CVPR’23) | 37.6 | 63.8 | 17.7 |
| Ours (Qwen-VL-2B) | 46.6 | 67.1 | 21.1 |
| Ours (Qwen-VL-7B) | 50.9 | 70.2 | 21.5 |
| Complete Baseline | -6.1 | -5.5 | - |
| SOTA | +7.6 | +4.3 | +3.8 |
I.2 Table of Ablation Study
We present complete results of the ablation study in Table 6.
| MM-IMDb | IU X-ray | |||||||||||||||||
| Missing Rate | 0.3 | 0.5 | 0.7 | 0.3 | 0.5 | 0.7 | ||||||||||||
| Variants | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS | F1 | AP | SS |
| Baseline (Qwen-VL-7B) | 54.7 | 60.9 | 33.5 | 54.9 | 61.3 | 32.7 | 55.2 | 61.8 | 32.3 | 53.6 | 73.9 | 22.6 | 50.9 | 70.2 | 21.5 | 46.3 | 70.5 | 19.8 |
| w/o Knowledge Modeling | -1.2 | -3.3 | -7.0 | -1.5 | -4.1 | -8.8 | -1.3 | -3.6 | -8.8 | -12.1 | -21.6 | -10.7 | -13.3 | 26.8 | -11.6 | -17.5 | -29.2 | -13.7 |
| + Random Ranking | -1.7 | -3.5 | -7.2 | -1.6 | -4.3 | -9.0 | -1.6 | -4.1 | -9.9 | -13.9 | -26.8 | -11.4 | -14.7 | -28.1 | -12.4 | -19.3 | -31.8 | -15.0 |
| Random Ranking | -0.5 | -2.6 | -0.6 | -0.4 | -2.8 | -0.7 | -0.5 | -2.7 | -0.6 | -2.9 | -6.3 | -3.8 | -3.1 | -6.9 | -4.3 | -3.8 | -7.1 | -4.7 |
| w/o Knowledge Ranking | -0.1 | -0.4 | -0.2 | -0.2 | -1.9 | -0.4 | -0.2 | -0.8 | -0.1 | -1.3 | -3.3 | -1.5 | -2.1 | -5.4 | -3.6 | -1.9 | -5.7 | -2.1 |
| w/o Semantic Ranking | -0.3 | -1.4 | -0.2 | -1.4 | -0.2 | -0.1 | -0.2 | -1.0 | -0.3 | -0.9 | -0.4 | -0.7 | -1.3 | -2.3 | -1.1 | -2.4 | -3.3 | -1.6 |
I.3 More Modalities Results
The proposed method primarily focuses on image and text modalities to facilitate the evaluation of the proposed method, as these modalities are well-supported by the community and computationally efficient when using large vision-language models. However, our approach imposes no constraints on the modality encoders, allowing it to be easily generalized to other modalities. To validate this, we conducted experiment on a multimodal sarcasm detection dataset [11], which includes audio, vision, and text modalities. Using Unified-IO [10] as the backbone and keeping other configurations unchanged, our method demonstrated effectiveness even when extended to these modalities, as shown in Table 7.
| Missing Rate | 0.3 | 0.5 | ||||
| Method | F1 | AP | SS | F1 | AP | SS |
| Baseline (complete) | F1: 62.4 mAP: 64.7 SS: - | |||||
| Baseline (remove missing) | 57.1 | 59.3 | - | 53.8 | 54.6 | - |
| DiCMoR (CVPR’23) | 55.5 | 57.6 | 11.7 | 51.3 | 52.9 | 10.2 |
| Ours (Unified-IO 7B) | 58.3 | 60.1 | 13.3 | 54.7 | 55.8 | 11.4 |
J Visualization Analysis
J.1 Completion Results
We present additional completion results, as shown in Fig. 4, 5, and 6. In the general domain, our knowledge modeling module emphasizes understanding the quantity of objects, their attributes, and the contextual environment. The results in Fig. 4 indicate that our method more closely resembles the original missing modality compared to direct generation approaches. In the medical domain, incorporating knowledge of different lesions allows the LMM to comprehend the relationships between various regions in chest X-rays and the content described by the modality. Figs. 5 and 6 demonstrate that our method offers a more reliable strategy for missing data completion than direct generation.
J.2 Intermediate Results
J.3 Knowledge-based Ranking Results
We present partial results of knowledge-based ranking as shown in Fig. 7. Here, “available [modality]” indicates that the modality is visible, and “QS” represents the quality score of the completed missing modality combined with the observed modality. The results demonstrate that our proposed knowledge-based ranking module can effectively select relatively reasonable generated outcomes.
K Hallucinations
The hallucinations of LMM [13] refer to instances when an AI model generates content that is factually incorrect, misleading, or unsubstantiated. In our approach, hallucinations primarily stem from the limitations of the training data. We advocate for a training-free method to achieve MMC, which has the advantage of being easily deployable across various domains with minimal input of relevant domain knowledge. However, the drawback is that the lack of task-specific training can result in the model having less “common sense” compared to models trained on extensive datasets. As illustrated in Fig. 10, the absence of common sense in our approach may lead to results that deviate from expected cognitive outcomes. In recent years, RAG (Retrieval-Augmented Generation) [27, 18] has been regarded as an effective technique for mitigating hallucinations in large models. This technique provides the model with external truths, thereby reducing hallucinations during the reasoning process. In future work, we plan to incorporate RAG to enhance the robustness of our approach.
L Limitations
L.1 More Modalities
Our method focuses exclusively on image and text modalities, leaving its performance on other modalities, such as speech and depth, yet to be explored. The approach emphasizes the automatic extraction of inter-modal knowledge and the completion of missing modalities through domain knowledge. However, this focus on a limited set of modalities limits its generalizability and adaptability in real-world applications where multi-modal data often involves various types of sensory inputs. Thus, in the future, adaptation to other modalities is possible by defining a more comprehensive modality knowledge and expanding the learning framework to accommodate these new modalities. Some promising works [14, 35] show that one modality, such as image or text, can be connected to any other modality, paving the way for more inclusive and versatile multi-modal systems that handle diverse data types with high efficacy.
L.2 More Tasks and Metrics
Additionally, we observe that while our method enhances classification performance under a high missing rate (e.g., 0.7), it paradoxically results in a decrease in the similarity scores of the completed modalities. This suggests that although the model performs well in reconstructing missing data for classification tasks, the semantic alignment and quality of the generated modalities may still require significant refinement. Addressing these limitations presents an opportunity to improve the balance between classification accuracy and modality similarity. Therefore, there remains substantial potential for further exploration to develop more robust generation and ranking strategies in the future. These improvements could include incorporating advanced similarity-preserving techniques and exploring diverse evaluation metrics to assess the completeness and coherence of generated data across different tasks.